OpenStudy (anonymous):

The measure of the supplement of an angle is 60 less than three times the measure the complement of the angle. Find the measure of the angle

OpenStudy (jack1):

supplement + complement = 180 or 360?

OpenStudy (anonymous):

suppl=180 compl=90

OpenStudy (anonymous):

127.5?

OpenStudy (jack1):

angle = x, compliment to x = c, supplement to x = s x+c = 90 x+s = 180 s = 3c - 60 solve for x using substitution

OpenStudy (jack1):

so: x+c = 90 ==> x = 90 - c so: x+s = 180 [90 - c] + s = 180 s = 180 - 90 + c s = 90 + c and s = 3c - 60 so: 3c - 60 = 90 + c 2c - 60 = 90 2c = 150 c = 75 so if c = 75 x + c = 90 x + 75 = 90 x = 15 so s = 175 c = 75 x = 15

OpenStudy (jack1):

cool, but did that all make sense dude? like, do u understand all the working out ok?

OpenStudy (anonymous):

yes thank you!!!

OpenStudy (jack1):

sweet, happy days yw, cya